OpenStudy (anonymous):

what is the slope of a line that is perpendicular to the lione represented by the equation 8x-2y=4 a)-1/4 b)-4 c)4 d)1/4

OpenStudy (anonymous):

8x-2y=4 -2y =4 - 8 x y = - 2 + 4x The slopes of lines perpendicular to one another when multiplied give -1. 4 * m = - 1 m = -1/4